The magnitude of result out of two equal forces is equal to either of two forces. The
angle between forces is
 

Grade:11

1 Answers

Aman
3 years ago
Let one force is F1
Second force is F2
Resultant force is F3
F1=F2=F3=F
NOW 
F3^2= F1^2 + F2^2 +2F1F2 Cos0
Cos0 = -1/2
0 = 120